You invest money into an account where the interest is compounded continuously. If the accounts value doubles in 9 years, what is the interest rate?
Answer by RAY100(1637)

continuous interest formula is A=Pe^rt,
,
A=Pe^rt
A/P= 2 = e^(r*9)
ln both sides
ln2 = 9r,,,,,,,,,,,,,,,,,,,,,,,,ln e^rt = rt
lnr/9 = .077 = r
,
check
A=Pe^rt
A/P = 2 = e^(.077 * 9) =2,,,,,ok